OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (anonymous):

@Chad123

OpenStudy (anonymous):

2 = 1*(1 + 1) = 2

OpenStudy (anonymous):

\(P_1\) is what you get when you replace \(n\) by \(1\) namely \[2=1(1+1)\]

OpenStudy (anonymous):

2 + 4 + ... + 2n + 2(n+1) = (n + 1)(n + 2)

OpenStudy (anonymous):

okay so i got that right for Pk I replaced the Ns with Ks. Is that right?

OpenStudy (anonymous):

\(P_k\) is what you get when you replace \(n\) by \(k\)

OpenStudy (anonymous):

\[2 + 4 + 6 + . . . + 2n = n(n+1) \]replace \(n\) by \(k\) get \[2 + 4 + 6 + . . . + 2k = k(k+1) \]

OpenStudy (anonymous):

Is that all i have to do to get Pk?

OpenStudy (anonymous):

yes believe it or not, that is all

OpenStudy (anonymous):

how do I do the end?

OpenStudy (anonymous):

replace \(n\) by \(k+1\) use parentheses

OpenStudy (anonymous):

so it would be k+1*(k+1)+1)?